On a trip to Kruger, you heat 1.5 kg of water on a camping stove. The water is originally at 23 °C, and you heat it long enough to boil 0.65 kg of it away. A. How much heat went into increasing the temperature of the water? Units: J

B. How much heat went into changing the phase of the water? Units: J

C. In your camping stove, 50% of the energy released in burning the fuel goes into heating the water. How much fuel did you burn in the above process? Given: 1 gram of fuel releases 46 kJ of energy.
Answer #1

Specific heat capacity of water, Cw = 4182 J/(Kg K)

Latent heat of vapourization of water, L = 22.6×105 J/Kg

Mass of water, Mw = 1.5 Kg

Mass of water converted to steam, MS = 0.65 Kg

A) Heat went into increasing temperature of water = Mw×Cw×(100-23) = 1.5×4182×77 = 483021 J

B) Heat went into changing the phase of water = L×MS = 22.6×105×0.65 = 1469000 J

C) Amount of fuel = 2×(1469000+483021)/46000 = 84.87 g = .08487 Kg